MORPHOLOGY OF GAS IN THE GALACTIC CENTER FROM SPECTROSCOPY OF H_{3}^{+}

\underline{TAKESHI OKA}, Department of Astronomy and Astrophysics and Department of Chemistry, University of Chicago, Chicago, IL 60637; THOMAS R. GEBALLE, Gemini Observatory, Hilo, HI 96720; NICK INDRIOLO, Department of Physics and Astronomy, Johns Hopkins University, Baltimore, MD 21218; MIWA GOTO, Universitäts-Sternwarte München, München Germany 81679.

Over the last several years our observations of the infrared spectrum of H_{3}^{+} toward the Galactic center (GC) have established a high ionization rate (ζ > 2 × 10^{−15} s^{−1}) and the existence of a vast amount of warm (T ∼ 250 K) and diffuse (n < 100 cm^{−3}) gas with a high volume filling factor (f > 0.3) in the Central Molecular Zone (CMZ) of the GC, a region of radius ∼ 150 pc.^{a}^{b}^{c} These findings are gradually being assimilated into the astrophysics of the GC.

Determining the morphology of this gas is difficult because the sightlines for study are limited by the uncontrollable locations of background stars suitable for spectroscopy of H_{3}^{+}. There are wide longitudinal gaps in the locations of those stars and the precise radial locations of the stars within the CMZ are uncertain. Nevertheless, the velocity profiles of the observed H_{3}^{+} spectra indicate the presence of the Expanding Molecular Ring (EMR), a structure containing mostly diffuse gas expanding from the center with velocities of up to 180 km s^{−1} and bordering the CMZ. On the other hand, the 120 pc Molecular Ring,^{d} an inner t ring of cold dust and dense gas with radius ∼ 100 pc is not clearly seen in H_{3}^{+}. This is possibly because the sightlines that we have observed to date lie close to the Galactic plane and miss the ring, which goes above and below the Galactic plane.^{e}

aOka, T., Geballe, T.R., Goto, M., Usuda, T., McCall, B.J. 2005, ApJ, 632, 882

bGoto, Usuda, Nagata, Geballe, McCall, Indriolo, Suto, Henning, Morong, and Oka, 2008, ApJ, 688, 306.

cGeballe, T.R. and Oka, T. 2010, ApJ, 709, L70

dSofue, Y. 1995, PASJ, 47, 527

eMolinari, S. et al. 2011, ApJ, 735, L33.